OpenStudy (anonymous):

convert to a percent 7 would 700% be correct?

Parth (parthkohli):

To convert a number into percentage, just get the denominator as 100 and then write the numerator(also attaching the percent sign)

OpenStudy (anonymous):

so my answer wasn't correct?

OpenStudy (asnaseer):

to convert a decimal number or a fraction to a percentage, you just need to multiply it by 100. e.g. 7 = 7 * 100 = 700% 0.5 = 0.5 * 100 = 50% 3/4 = 3/4 * 100 = 75% you had asked the same question before: http://openstudy.com/updates/4fcd40a7e4b0c6963ad8374e